### v1

# BillBoard Splatting (BBSplat): Learnable Textured Primitives for Novel View Synthesis

Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. Our method’s qualitative and quantitative improvements over 3D and 2D Gaussians are most noticeable when fewer primitives are used, when BBSplat achieves over 1200 FPS. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. We demonstrate improvements on PSNR, SSIM, and LPIPS metrics compared to the state-of-the-art, especially for the case when fewer primitives are used, which, on the other hand, leads to up to inference speed improvement for the same rendering quality.
1 Introduction
Novel view synthesis (NVS) is a crucial technology for a variety of applications, including virtual reality, computer gaming, and cinematography. Several efforts have been dedicated to deploy methods that are more efficient while providing a better quality of the synthesized images. Specifically, the choice of geometric primitives used to represent the scene plays a key role in defining the advantages and drawbacks of different NVS methods. Recent breakthroughs on neural representations [30, 31, 34, 4, 5, 35, 47, 8, 18, 7, 51] have been sided by recent methods based on Gaussian Splatting [19, 23, 24, 52, 10, 36, 13, 54], demonstrating most efficient way for novel view rendering.
Indeed, NeRF-based methods [34, 5, 4, 35] still achieve the best NVS quality for challenging real-world captures by using implicit scene representations such as the weights of an MLP. However, image rendering with a NeRF is less efficient as it requires repeated MLP inferences to predict colors along camera rays. An alternative to neural rendering, 3D Gaussian Splatting (3DGS) [23] used faster rendering based on projecting explicit primitives on the screen surface while preserving high-quality NVS. In practice, 3DGS uses Gaussian-distributed radiance around explicit 3D scene points as primitives.
Recently, 2D Gaussian Splatting (2DGS) [19] proposed the use of flat Gaussians, oriented in 3D, to represent the scene more efficiently. Since 2D Gaussians are effectively tangent to object surfaces, they allow more accurate surface extraction. Despite proving their efficiency in mesh extraction task, 2D primitives result in a downgrade of rendering metrics compared to 3D primitives such as 3D Gaussians. In this work, we aim to make 2D primitives suitable for high-quality NVS by intoducing a new primitive representation.
Our proposed geometric primitives for NVS take inspiration from the classic billboards used for extreme 3D model simplification [9] by replacing mesh during a greedy optimization process. A 3D scene can be efficiently rendered by using a “billboard cloud” of given textured planar primitives with alpha channels. Using billboards, we can efficiently model planar surfaces, such as a painting on a wall or scene background, while dramatically reducing the number of geometric primitives required, up to an order of magnitude with respect to 3DGS/2DGS (check LABEL:fig:shapes for a comparison). In combination with efficient texture sampling implemented on the GPU, this leads to inference speed improvements without a drop in the rendering quality (Section 4.2).
We define billboards with 2D Gaussians parameters (rotation, scaling, 3D center location, and spherical harmonics) while also introducing RGB texture and alpha map to control pixel-wise color and shape (Fig. 1(b2)). The alpha map defines the billboard silhouette and models the arbitrary shape of primitives. Similarly, the RGB texture stores color for each point of the billboard. In this way, we can use fewer primitives to represent high-frequency details (LABEL:fig:shapes). The key aspect of our approach, BillBoard Splatting (BBSplat), is a method to learn billboard parameters from a set of calibrated images.
To ensure rendering efficiency, we implemented texture sampling and a back-propagation process in CUDA. In this way, the use of textures does not result in overtime compared to the 2DGS Gaussians’ rasterization process. To tackle the challenge of storing all billboard textures, we compress them by representing each texture as sparse offsets from colors predicted by spherical harmonics [2] and by further quantizing them to 8 bits. Then, we can efficiently utilize dictionary-based compression algorithms [41, 11] for quantized textures.
To summarise, our contributions are as follows:
We propose BBSplat with optimizable textured primitives to learn 3D scene representation with photometric losses. BBSplat allows up to acceleration for NVS compared to the state-of-the-art for the same rendering quality.
We developed an algorithm to efficiently represent and store textures for billboards. BBSplat demonstrates significant storage cost reduction. As a result, for some scenarios, storing the scene represented with billboards is more efficient than 3DGS or 2DGS.
We implemented BBSplat with CUDA and demonstrated its efficiency in extensive experiments on several real-scene open datasets.
2 Related work
Reconstruction of 3D scene for novel-view synthesis is a long-standing problem that used to be solved with structure from motion approaches (SfM) [40, 56, 45, 15, 44]. The significant increase in quality was achieved with the appearance of implicit scene representation [30, 34, 37].
Nowadays, Neural Radiance Fields (NeRF) [34] is the most commonly used implicit method for NVS. It uses a neural network to predict points’ colors and transparency based on points coordinates along the camera ray. The biggest drawback of such representation is rendering speed, as it requires multiple inferences of an MLP. InstantNGP [35] achieves the most acceleration of NeRF up to 150 FPS by using multi-scale feature grids.
To overcome the rendering speed issue, the 3D Gaussian Splatting (3DGS) [23] method proposed an explicit method for 3D scene representation. Based on the point cloud reconstructed by SfM [40], they train the color and orientation of ellipsoids with Gaussian distributed transparency. For NVS, these ellipsoids are splatted on the screen surface providing high rendering speed. As a result, 3DGS has become popular for many practical real-time applications such as human avatars [46, 29, 50] or SLAM [32, 53].
Recent works further improve 3DGS in terms of rendering quality and speed. Thus, in [6] authors revise the densification algorithm to reduce the number of artifacts by using loss-based splitting of Gaussians. In 3DGS-MCMC [24], authors use Markov Chain Monte Carlo interpretation of sampling algorithm to more efficiently distribute Gaussians over the scene, even with random initialization. RadSplat [36] uses pre-trained NeRF representation as a prior during 3DGS training to improve rendering quality in challenging scenarios. Mip-Splatting [52] solves the issue of gaps between Gaussian produced due to changes in camera distance by introducing a 3D smoothing filter that constrains Gaussian size.
Some of the recent works also propose to disentangle appearance and geometry for 3DGS. Texture-GS [49] introduces an MLP to establish correspondence between 3D Gaussian surface and texture map. In practice this approach allows one to edit the texture in the scene but decrease rendering speed to three times. Another approach [20] encodes color and opacity as spherical harmonics on the surface of 3D Gaussian, which allows one to improve reconstruction fidelity.
2D Gaussian splatting (2DGS) [19] was proposed to increase the quality of mesh surface reconstruction from such scene representation. Specifically, 2DGS uses ellipses instead of ellipsoids as they align along the object’s surface and make mesh extraction operation more straightforward. For faster rendering of such flat primitives, 2DGS utilizes a ray-splat intersection algorithm [48, 42]. While 2DGS increases the accuracy of mesh reconstruction, its fidelity falls shortly behind 3DGS, especially for in-the-wild scenes. In this work, we investigate the ability of flat primitives to be used for high-quality novel-view rendering.
We noted that existing approaches use a tremendous number of Gaussians to represent the scene, which affects rendering speed. Billboard clouds [9] propose an extreme 3D model simplification by replacing it with a set of textured planes. This representation is widely used in computer graphics for real-time forest rendering [3, 14, 27] and found implementation in such tasks as automatic paper slice-form design [28, 33, 39] and LOD control [17]. In this work, we take inspiration from the billboard clouds technique [9] and propose arbitrarily shaped textured primitives as an extension of the 2DGS approach, resulting in faster and more qualitative rendering of real scenes reconstructed from the set of photographs.
3 Method
3.1 Preliminaries: 2D Gaussian Splatting
Recently proposed 2D Gaussian splitting (2DGS) [19] uses 2D flat ellipses (in contrast with 3D ellipsoids for 3DGS) orientated along the object’s surface to represent the scene. These ellipses have Gaussian distributed transparency and are parametrized as which corresponds to position, 2D scale, quaternion rotation, global opacity, and spherical harmonics of the primitive as in Fig. 1(b1). Here spherical harmonics [2, 12] define view-dependent color. In practice, to calculate ellipses orientation, 2DGS uses local tangent planes defined with transformation matrix as:
where matrices and produced from quaternion rotation and scale , and is the point coordinate on the plane . Then, 2DGS utilizes coordinates of points within a plane to calculate 2D Gaussian intensity as follows:
To project Gaussians on the screen, 2DGS finds coordinates corresponding to the screen pixel as the intersection of a camera-direction ray with a plane . For efficiency, 2DGS utilizes an explicit ray-splat intersection algorithm proposed by [42].
3.2 Billboard Splatting
In this work, we propose to further improve NVS efficiency and quality by introducing billboard splatting. These billboard primitives are based on planes, but leverage learnable textures instead of utilizing Gaussian distribution (Eq. 2) to calculate their color and transparency. We inherit 2DGS parametrization for these planes: , where is the center of -th plane, is scales along two axis, is rotation quaternion, is Spherical Harmonics coefficients. Instead of using global Gaussian opacity we set transparency at each point of a plane as texture (Fig. 1 (b2)). This way, each billboard can have an arbitrary shape. Additionally, we parametrize billboards with - RGB texture to control the color of all points of the billboard.
We use explicit ray-splat intersection algorithm [42] to find the corresponding plane coordinate for screen coordinate. Specifically, we parametrize the camera ray as an intersection of two 4D homogeneous planes: and . We then transform these planes to the -coordinate system of plane to find intersection point in coordinates:
where is the transformation matrix from world space to screen space. Then we calculate the coordinate of the intersection point as:
where are the -th value of the 4D homogeneous plane parameters.
This way, we get coordinates in the range with corresponding to the plane center . After that, we rescale them to the range , where corresponds to texture size in texels. Using rescaled coordinates, we sample color and opacity of the ray-splat intersection point (Fig. 1 (c2)) and accumulate them along the ray as:
where is the view-dependent color calculated with and sampled color . To sample textures’ values , we utilize bilinear sampling [43]. This way, we take into account neighboring texels and can calculate gradients for billboard position with respect to the texture. We adopt PyTorch [38] CUDA implementation of bilinear sampling to sample texture values and calculate gradients. Namely, we redistribute input gradient values between texels that contributed to the sampled value according to their contribution weights (i.e. bilinear coefficients).
To calculate value of view-dependent colors at each uv-point u of -th billboard, we integrate sampled texture colors as offsets to the colors predicted with and view-direction vector :
Here is a spherical harmonic features [2] for each primitive, and is the view-direction vector from the camera position to billboard center position . We transform to RGB color space and define this operation as . This way, we can handle view-dependent light effects for textured planes (Fig. 2). Representing the final color as a combination of spherical harmonics and sampled color was initially proposed in [49] and helps to disentangle texture and light effects. In our case it also allows to have a sparse structure, which is more efficient to store as described in Section 3.5.
Final parametrization of billboard is as follows: , where follows 2DGS parametrization (Section 3.1) and the values correspond to our proposed textures for color and opacity.
3.3 Training
To train billboard splatting representation of the scene, we initialize them with sparse point cloud obtained using SfM [40]. Optionally, we add 10K evenly distributed points on the circumscribing sphere using the Fibonacci algorithm [16] to represent the sky and far away objects. We use only photometric losses to fit a scene as in Gaussian splatting pipelines. For textures, we propose regularizations to avoid their overfitting and get a more sparse structure that reduces storage costs. In this section, we describe the losses and regularizations we used.
Photometric losses. We utilize photometric losses to train our billboard representation. Namely we apply and structure similarity losses:
Therefore, we need only a set of images to train the BBSplat representation of the scene.
Regularizations. Due to the large number of parameters, textures tend to overfit training images, which results in noisy rendering output for novel views. To alleviate this issue, we propose a simple yet efficient regularization in which we push billboards with low impact on the renders (i.e. small or far billboards affecting few pixels) to have Gaussian distributed transparency.
First, we define per-texture visibility weight for all billboards based on their impact on the rendered image. We define the impact for -th billboard as a sum over alpha-blending values corresponding to its rendered pixels :
Alpha-blending values are calculated similarly to 2DGS and take into account overlapping planes at pixel . We use the maximum impact threshold to regularize only clearly visible billboards. The choice of the threshold is discussed in the supplementary materials. By ignoring billboards with , we do not regularize invisible in the frame primitives, avoiding over-regularization in underrepresented areas. The proposed criterion is more efficient than the one utilizing projected radii as in 2DGS since they only take hitting frustum into account and ignore overlapping.
Finally, we define the regularization term to push RGB textures close to zero and alpha-maps close to a Gaussian distribution based on their visibility weight :
With this regularization, the final loss to train BBSplat 3D scene representation is defined as:
3.4 Adaptive density control
Most 3DGS-based methods use adaptive density control to adjust the number of Gaussians. Specifically, they perform splitting, cloning, and pruning operations according to user-set policies. An alternative approach is proposed in [24], where they reformulate Gaussian Splatting densification as a Markov Chain Monte Carlo (MCMC) sampling. MCMC introduces noise in the training process and replaces splitting and cloning with a deterministic state transition.
We also adopt and modify the MCMC sampling technique for BBSplat. In the sampling stage, MCMC clones Gaussians and adjusts their opacity and scale set by covariation matrix to preserve the rendering output state:
While this adjustment is well fit for Gaussians, we noticed that scale adjustment in Eq. 12 is unsuitable for arbitrarily shaped primitives and violates the rendering state preservation. Instead, we adjust only to preserve overall transparency along the ray, because it controls transparency in all points on a plane and therefore does not require additional scale modification:
This renders parameters adjustment more straightforward compared to Eq. 12. Finally, where MCMC compares opacity with a predefined threshold to determine “dead” Gaussians, we compare the average value of the alpha-map: .
3.5 Texture compression
Using textured splats leads to new challenges of efficiently storing them. For each point, we use two textures of size with one channel for transparency and three channels for color. Considering that textures are stored with 4-byte float point values, we need bytes of memory to store each texture.
To reduce storage costs, we first apply quantization [21] of texture parameters. Let us define normalized in range texture values as . For storing BBSplat, we quantize texture values by rescaling where is the rounding operation. When loading textures, we apply an inverted operation for dequantization. This way, we can store each texel value as a single byte, reducing storage costs four times.
Our final color representation described in Eq. 6 as a sum of color calculated with SH and texture sampled offsets allows us to additionally reduce storage costs. Our regularization term proposed in Eq. 9 allows us to get sparser values and closer to a Gaussian distribution . We increase the sparsity of by subtracting the Gaussian pattern and ensure more zero values in the texture. The sparse nature of textures allows us to use efficient dictionary-based compression methods [41, 11] (e.g. ZIP [1]) to further reduce storage costs. Our texture compression pipeline allows us to reduce memory consumption times on average.
4 Experiments
We compared our billboard splatting representation against five state-of-the-art 3D scene representation methods: 3DGS [23], Mip-Splatting [52], 3DGS-MCMC [24], Mini-Splatting [10], and 2DGS [19]. For 2DGS, we conducted two sets of experiments: with and without depth-normal regularizations, as they affect metrics differently for various datasets. In the following section, we provide both quantitative and qualitative comparisons for indoor and outdoor real scenes from standard datasets: Tanks&Temples [26], Mip-NeRF-360 [5], and DTU [22]. Please check supplementary material for additional images, quantitative results, and further details on experimental implementation and evaluation.
Tanks&Temples (outdoor) DTU Mip-Nerf-360 (indoor) Points FPS Storage PSNR SSIM LPIPS Points FPS Storage PSNR SSIM LPIPS Points FPS Storage PSNR SSIM LPIPS 3DGS [23] 80K 1251 ± 187 19 MB 22.25 0.8185 0.2715 30K 1483 ± 228 7 MB 27.28 0.9115 0.2022 160K 1374 ± 127 38 MB 27.56 0.8978 0.1867 Mip-Splatting [52] 80K 1162 ± 180 19 MB 22.29 0.8198 0.2711 30K 1355 ± 195 7 MB 26.89 0.9072 0.2045 160K 1254 ± 104 38 MB 27.25 0.8948 0.1899 3DGS-MCMC [24] 80K 1158 ± 106 19 MB 25.07 0.8619 0.2276 30K 1186 ± 181 7 MB 30.96 0.9365 0.1741 160K 1109 ± 110 38 MB 30.71 0.9264 0.1417 Mini-Splatting [10] 80K 1113 ± 152 19 MB 24.18 0.8358 0.2519 30K 1151 ± 188 7 MB 29.21 0.9236 0.1836 160K 1143 ± 102 38 MB 30.23 0.9178 0.1437 2DGS [19] 80K 1269 ± 113 18 MB 17.83 0.6918 0.4222 30K 1258 ± 191 7 MB 27.75 0.9105 0.2039 160K 1174 ± 83 37 MB 26.32 0.8755 0.2177 2DGS† [19] 80K 1292 ± 105 18 MB 22.07 0.8154 0.2793 30K 1249 ± 187 7 MB 28.69 0.9184 0.1965 160K 1207 ± 84 37 MB 27.96 0.8971 0.1888 Ours 80K 1217 ± 92 49 MB 25.52 0.8782 0.1708 30K 1197 ± 181 15 MB 30.14 0.9387 0.1472 160K 1093 ± 76 98 MB 31.33 0.9356 0.1078
Tanks&Temples (outdoor) DTU Mip-Nerf-360 (indoor) Points FPS Storage PSNR SSIM LPIPS Points FPS Storage PSNR SSIM LPIPS Points FPS Storage PSNR SSIM LPIPS 3DGS [23] 830K 724 ± 75 196 MB 24.77 0.8752 0.1828 300K 964 ± 135 71 MB 29.18 0.9335 0.1573 1233K 599 ± 35 292 MB 31.63 0.9369 0.1144 Mip-Splatting [52] 1114K 596 ± 49 268 MB 24.66 0.8813 0.1714 431K 787 ± 95 104 MB 29.39 0.9333 0.1362 1562K 467 ± 19 375MB 31.72 0.9405 0.1025 3DGS-MCMC [24] 830K 680 ± 35 196 MB 26.19 0.8940 0.1639 300K 833 ± 133 71 MB 31.41 0.9511 0.1143 1233K 519 ± 34 292 MB 32.30 0.9448 0.1013 Mini-Splatting [10] 193K 992 ± 159 46 MB 24.62 0.8782 0.1776 215K 852 ± 142 51 MB 27.74 0.9257 0.1301 293K 1001 ± 78 70 MB 30.97 0.9352 0.1105 2DGS [19] 460K 947 ± 80 107 MB 21.75 0.7807 0.3091 151K 1101 ± 194 35 MB 29.83 0.9339 0.1628 764K 715 ± 46 163 MB 30.26 0.9178 0.1519 2DGS† [19] 542K 851 ± 72 126 MB 24.95 0.8736 0.1912 135K 1109 ± 193 31 MB 29.85 0.9343 0.1612 674K 763 ± 51 157 MB 31.37 0.9319 0.1243 Ours-small 160K 1044 ± 89 93 MB 25.68 0.8797 0.1649 30K 1197 ± 181 15 MB 30.14 0.9387 0.1472 160K 1093 ± 76 98 MB 31.33 0.9356 0.1078 Ours-big 300K 900 ± 95 163 MB 25.78 0.8804 0.1625 60K 1178 ± 180 29 MB 30.10 0.9371 0.1449 300K 920 ± 69 173 MB 31.51 0.9371 0.1044
| PSNR | L1 | LPIPS | FPS | |
| Texture-GS | 30.03 | 0.0135 | 0.1440 | 267 |
| Ours-small | 30.81 | 0.0123 | 0.1055 | 1230 |
4.1 Implementation
We use for the structure similarity in the Equation 7, and for the Equation 9. We initialize and with zeros and 2D Gaussian transparency respectively, and keep them fixed for the first 500 iterations to adjust orientations and colors initialized with SfM. The densification strategy described in Sec. 3.4 is used from 500 until 25,000 iterations to adjust the number of billboards. Overall, we train BBSplat for 30,000 steps and finally, we fine-tune the spherical harmonics for an additional 2,000 iterations to adjust them for textures.
For training, we use the Adam [25] optimizer with hyper-parameters recommended for the Gaussian splitting pipelines [23]. We set RGB and textures size . resulting in textures to get the best trade-off between quality and memory requirements. To optimize RGB-texture we set learning rate and for -texture we set . Additionally, we found it beneficial for convergence to increase spherical harmonics learning rate to .
We implemented texture sampling with back-propagation in the CUDA code to achieve fast inference.
4.2 Results and Evaluation
We compared with previous state-of-the-art methods on real-world scenes using standart datasets. We evaluated our method on 4 indoor scenes from the Mip-NeRF-360 dataset [5] and 5 outdoor scenes from the Tanks&Temples dataset [26]. Additionally, we evaluated our method on 15 scenes from DTU dataset [22]. Overall, we provide experiments for 24 real scenes from three datasets.
We used train/test split for all datasets following 3DGS methodology where each 8th image was taken for the test. For all three datasets, we used images downsampled to 800 pixels on the larger side. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
Quantitative comparison. We report PSNR, SSIM, and LPIPS [55] metrics (Tabs. 1 and 2) widely used to evaluate quality in the NVS task. PSNR and SSIM are traditionally used to demonstrate similarity with ground truth images; LPIPS utilize neural networks and is known for the best correlation with human perception. To assess model efficiency, we also report rendering speed as frames per second (FPS) and required storage space along with a number of used primitives. In the provided experiments, we evaluated metrics for different numbers of primitives by limiting the growth of their number during training as proposed in [6].
In an experiment in Table 1, we showcase metrics for the fixed number of primitives depending only on their type. For the same amount of primitives, our method reaches better PSNR, SSIM and LPIPS than state-of-the-art Gaussian-based representations both for indoor and outdoor scenes. While improving state-of-the-art objective metrics, for the same number of primitives BBSplat requires more storage space. In the next experiment, we demonstrate inference speed and storage efficiency for on-par metrics.
In Table 2, we compare against other methods for the maximum recommended amount of Gaussians to achieve the best possible quality for each method. For our method, we provide two scenarios with fewer (small) and more (big) billboards. We outperform 3DGS and 2DGS for all three metrics, at the same time showing inference speed and storage space improvement. BBSplat demonstrates metrics on par with the best competitor, 3DGS-MCMC, achieving up to two times faster rendering of the scenes with three times lower storage space requirement.
In Table 3, we compare with Texture-GS [49] as it also proposes to employ textures, although in conjunction with a 3DGS model. The main objective of Texture-GS is texture editing: indeed we verify that Texture-GS performance are reduced for the NVS task, especially in FPS count. We report metrics only for the DTU dataset since Texture-GS is unable to process large scenes, as it is limited by spherical texture space definition.
In Figure 3, we show the metrics as a function of the number of primitives (Gaussians or BBSplats) and as a function of FPS. To plot the leftmost point we used the maximum number of Gaussians requested by each method, then for subsequent points we fixed the number of primitives for all scenes at the chosen value. It can be seen that our method gives a flatter curve of metrics change, which means a minimal decrease in metrics when the number of primitives decreases to 80K. Thus, the proposed implementation allows us to achieve an inference speed of up to 1250 FPS while maintaining high rendering quality.
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art while providing faster inference speed and higher level of detail. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to supplementary materials.
4.3 Ablation study.
| PSNR | SSIM | LPIPS | Storage | |
| w/o | 25.17 | 0.8684 | 0.1774 | 60 MB |
| w/o | 25.26 | 0.8710 | 0.1858 | 72 MB |
| w/o | 25.54 | 0.8791 | 0.1655 | 131 MB |
| w/o compress | 25.71 | 0.8818 | 0.1630 | 662 MB |
| Full | 25.68 | 0.8797 | 0.1649 | 93 MB |
In Table 4, we provide an ablation study of different aspects of our approach. Namely the use of RGB and alpha textures, texture regularization term, and compression post-processing.
The use of both alpha and RGB textures is essential for rendering quality. Excluding significantly increases LPIPS, while excluding reduces PSNR and SSIM. Texture regularization term preserves over-fitting and leads to additional metrics improvement. It also encourages sparsity in textures and reduces storage costs in combination with compression. The use of compression leads to a slight metric reduction but allows to significantly reduce storage space requirements.
Therefore, all parts of the proposed approach are essential to get the best result. However, we note that metrics could be additionally improved by disabling compression.
5 Conclusion
We have proposed BBSplat - a novel method for 3D scene representation. The proposed method uses textured primitives of arbitrary learnable shapes and allows a two times faster generation of photorealistic novel views. We developed specialized regularization term and compression technique to reduce storage space by leveraging the sparse nature of the proposed textured representation. In extensive experiments on real data, we demonstrated the efficiency of the proposed method both quantitatively and qualitatively. In particular, we showed the best trade-off between objective metrics and inference speed.
Limitations and future work. One of the main limitations of the proposed method is storage space. While we make a significant step towards its reduction, there is still room for improvement in future works. Another limitation is training time, as it takes about 40 minutes to fit a scene, compared with 5 minutes for 3DGS. This slowdown is caused by backpropagation to the textures, and while it still significantly outperforms NeRF-based methods in terms of speed, it could be a bottleneck for some applications. In future works, we are going to focus on resolving these limitations.
Acknowledgements.
We thank Matteo Toso, Milind Gajanan Padalkar, Matteo Bortolon, and Fabio Poiesi for their feedback and helpful suggestions on improving the paper. We also thank Andrea Tagliasacchi for insightful discussions and Matteo Bortolon for the help with CUDA implementation. D. Svitov thanks the ELLIS network for organizing the PhD program that made this work possible.